seema wants 5 liter milk shopkepper has only 500ml how many packs of 500ml will she have to buy to get 5 liter milk
Answer by stanbon(75887)

seema wants 5 liter milk shopkepper has only 500ml how many packs of 500ml will she have to buy to get 5 liter milk
-----
Note: 500ml = (500/1000)liter = (1/2) liter
-----
10*(1/2)liter = 5 liters
-------
She needs to buy ten 500 ml packs
==================
Cheers,
Stan H.